0 0
Add a comment Improve this question Transcribed image text
Answer #1

a) Determine the best possible payoff, and choose the alternative with that payoff.

OK = 150

TX = 140

LA = 130

Ans - OK

b) Determine the worst possible payoff for each alternative, and choose the
alternative that has the “best worst.”

OK = 85

TX = 90

LA = 110

Ans - LA

c) Determine the worst regret for each alternative, and choose the alternative with the “best worst.” To do
this, subtract every payoff in each column from the best payoff in that column. The second step is to identify the worst regret for each alternative

Low Med High
OK 25 10 0
TX 20 20 10
LA 0 0 20

OK - 25

TX - 20

LA - 20

Ans - TX & LA either can be chosen

d) Minimize cost - It will be basis the weighted average of demand and value for each option:

OK = 85*20% + 110*30% + 150*50% = 125

TX = 90*20% + 100*30% + 140*50% = 118

LA = 110*20% + 120*30% +130*50% = 123

Hence, OK is the best